A car dealer advertised a big sale by stretching a string of banners from the top of the building to the edge of the driveway. If the building is 26 m high and the driveway is 39 m from the building,how long is the string of banners?
Answer by Boreal(15235)

The banner length is the hypotenuse
so 26^2+39^2=x^2=2197
x=46.87 m
this is a 2-3-sqrt 13 right triangle if the legs are divided by 13.
The hypotenuse is sqrt (13), and when scaled up, it is 13 sqrt (13), which is the answer above.